# From Personas to Talks: Revisiting the Impact of Personas on LLM-Synthesized Emotional Support Conversations

### v1

@@ Abstract
Shenghan Wu1*       Yang Deng2       Yimo Zhu1       Wynne Hsu1       Mong Li Lee1 1National University of Singapore             2Singapore Management University shenghan@nus.edu.sg, ydeng@smu.edu.sg, yimo@u.nus.edu, {whsu, leeml}@comp.nus.edu.sg
@@ 1 Introduction
In emotional support conversations (ESC), the supporter aims to help the seeker to reduce stress, overcome emotional issues, and promote mental well-being. Traditionally, ESC corpora have been developed through skilled crowdsourcing Rashkin et al. (2019); Liu et al. (2021), transcription of therapist sessions Liu et al. (2023); Shen et al. (2020), or by compiling emotional question-answer pairs from online platforms Sharma et al. (2020a); Sun et al. (2021); Garg et al. (2022); Lahnala et al. (2021). However, beyond the high costs, recent research Deng et al. (2024) highlights several limitations in these methods, including privacy concerns, variability in data quality, and fabricated user needs created by crowdworkers. With the advent of large language models (LLMs), their powerful generalization abilities enable high-quality data annotation and generation based on specific instructions Tan et al. (2024); Ding et al. (2024). Consequently, more and more recent studies Zheng et al. (2023); zheng2024self; Qiu and Lan (2024); Wu et al. (2024) investigate the use of LLMs to generate large-scale emotional support dialogue datasets across various scenarios via role-playing at lower costs. These efforts have significantly expanded the available ESC corpora.
Despite the significant strides of LLMs in synthesizing ESC, a significant issue in generative AI annotations is the Lack of Human Intuition Deng et al. (2024). Recent studies have demonstrated that effective emotional support requires careful consideration of individual differences, including personality traits, emotional states, and contextual factors (Ma et al., 2024; Ait Baha et al., 2023; Hernandez et al., 2023). This understanding has led to increased research attention on the role of persona in emotional support dialogues (Zhao et al., 2024; Cheng et al., 2022), particularly as AI systems become more prevalent in providing such support. The importance of incorporating psychological perspectives in developing empathetic AI assistants has been emphasized by researchers. Huang et al. (2023b) argues that psychological analysis of LLMs is crucial for creating more human-like and engaging interactions. While recent work has made progress in measuring LLMs’ personality characteristics using established psychological inventories (Frisch and Giulianelli, 2024; Safdari et al., 2023), there remains a gap in understanding how these persona-related aspects influence the generation of emotional support dialogues.
@@ 2.2 Emotional Support Dialogues
Early efforts on emotional support dialogues focused on collecting emotional question-answer data from online platforms (Medeiros and Bosse, 2018; Sharma et al., 2020b; Turcan and McKeown, 2019; Garg et al., 2022). These datasets laid the groundwork for understanding user emotions, but were limited to single-turn interactions. Empathetic Dialogue dataset (Rashkin et al., 2019) addressed this by introducing multi-turn dialogues, crowd-sourced to simulate diverse empathetic interactions. ESConv (Liu et al., 2021) further advanced the field by introducing emotional support strategies collected from psychological theories, enabling chatbots to use these strategies for more empathetic and contextually appropriate responses. Subsequent studies proposed using hierarchical graph networks to capture global emotions causes and user intentions (Peng et al., 2022), combining multiple emotional support strategies to enhance empathy (Tu et al., 2022), and implementing emotional support strategies and scenarios using LLMs to create the ExTES dataset (zheng2024self).
@@ 2.3 Persona-Driven Emotional Support
Recent advances have integrated personas into emotional support dialogues to enhance personalization and diversity. The ESC dataset (Zhang et al., 2024) introduced personas into the dialogue generation process. (Zhao et al., 2024) proposed a framework to extract personas from existing datasets for evaluation. Additionally, personas have been incorporated into chatbots to generate personalized responses (Tu et al., 2023; Ait Baha et al., 2023; Ma et al., 2024). These developments inspire our analysis of the relationship between personas, emotional support strategies, and dialogues, focusing on the extraction and use of personas in ESC.
@@ 3 Dataset Collection
In order to study the relationship between LLM-generated emotional support conversations (ESC) and persona traits, we first need to collect a set of real ESC data with user personas as reference. Specifically, we select three existing emotional support datasets, including ESConv (Liu et al., 2021), CAMS (Garg et al., 2022), and Dreaddit (Turcan and McKeown, 2019). ESConv is a multi-turn emotional support dialogue dataset, while CAMS and Dreaddit are derived from Reddit posts discussing mental health issues.
@@ 4 Measurement of Persona Traits (RQ1)
(a) Correlations between CSI and HEXACO from ESConv measured by gpt-4o-mini. P_value of all metrics are less than 0.01.
@@
(b) Correlations between CSI and HEXACO from ESConv measured by Claude-3.5-haiku.
@@
(c) Correlations between CSI and HEXACO from ESConv measured by LLaMA-3.1-8B-Instruct.
@@ 4.2 Results and Discussions
Based on the responses from these inventories, we calculate the HEXACO and CSI dimension scores for each dataset. To evaluate whether LLMs can infer stable traits from persona cards in an emotional support context, we then compute the correlations between the HEXACO and CSI dimensions within each dataset. Each scale provides a score corresponding to each response, along with the dimension to which each question belongs. Based on the LLM’s responses, we calculate the scores for each dimension. Finally, we use Pearson correlation to analyze the relationships between each dimension of HEXACO and CSI. Tables 2(a), 2(b), and 2(c) present the correlations between HEXACO and CSI dimensions in the ESConv dataset (Liu et al., 2021) as measured by three different LLMs. Similarly, Tables 6(a), 6(b), and 6(c) show the correlations in the CAMS dataset (Garg et al., 2022), while Tables 7(a), 7(b), and 7(c) provide the correlations observed in the Dreaddit dataset (Turcan and McKeown, 2019).
@@ 5.2 Results and Discussions
The comparisons of scores of persona traits are shown in the Figure 7, 7, and 7. The results show that the original personas and the personas extracted from the generated dialogues share similar distributions across four personality traits: Honesty-Humility, Agreeableness, Conscientiousness, and Openness to Experience. This indicates that the personas maintain consistent traits in the synthetic emotional support dialogues. However, we notice that the extracted personas tend to have higher Emotionality and lower Extraversion compared to the original personas. We believe this may be because the seeker in the emotional support dialogues is dealing with emotional issues, making them more emotional and less outgoing. A similar pattern is observed in the comparison of CSI scores (see Table 21, 21, and 21 in Appendix E). This indicates that the persona traits general maintain consistent in the synthetic emotional support dialogues with only slight shifts.
@@ 6 Impact of Persona on LLM-simulated Emotional Support Dialogues (RQ3)
Many efforts have been made on exploring how to incorporate emotional support strategies into emotional support dialogues. zheng2024self use LLMs to introduce emotional support strategies in synthetic dialogues. Zhang et al. (2024) further involve the concept of persona into these dialogues and analyze the usage of strategies. This raises an important question: does adding a persona have an impact on the way emotional support strategies (definitions of each strategy are shown in Appendix G) are distributed in dialogues? To investigate it, we use ESConv dialogues as history and instruct LLMs to predict how a new conversation might unfold after some time. Each future dialogue is generated in two versions: with and without persona traits (personality and communication style scores). Since ESConv provides real-world context and continuations generated by LLMs can preserve the semantic distribution of dialogues (Fan et al., 2025), our approach ensures diversity and realism in emotional support scenarios.
@@
(a) Strategy distribution on two different groups of synthesized dialogues by gpt-4o-mini, one generated with persona traits (PT), the other without.
@@
(b) Strategy distribution on two different groups of synthesized dialogues by claude-haiku-3.5.
@@
(c) Strategy distribution on two different groups of synthesized dialogues by LLaMA-3.1-8B-Instruct.
@@ 6.1 Analysis of Emotional Support Strategies
As demonstrated in Tables 3(a) and 3(b), the distribution of emotional support strategies differs significantly between dialogues generated with and without persona traits for both gpt-4o-mini and claude-3.5-haiku. While Table 3(c) shows that llama-3.1-8B-instruct exhibits less pronounced differences in strategy distributions between persona and non-persona conditions, the trends remain consistent with those observed in other models. The analysis revealed that supporters in persona-enhanced dialogues demonstrated a greater tendency to validate seekers’ emotions through questioning and provided more effective emotional comfort, as evidenced by significantly higher rates of questioning, affirmation, and reassurance strategies compared to non-persona dialogues. Conversely, supporters in dialogues without persona traits emphasized problem explanation and relied more heavily on self-disclosure, employing higher proportions of restatement, paraphrasing, and sharing of personal experiences as empathy-building mechanisms. The distributions of emotional support strategies generated using either personality scores or communication style scores, as shown in Table 4, demonstrate remarkable similarity. This alignment can be attributed to the strong correlation between these trait measures, as discussed in Section 4. These results provide compelling evidence that persona traits significantly influence the distribution and application of emotional support strategies in dialogues.
@@ 6.2 Human Evaluation
To intuitively reveal the impact of persona on LLM-simulated emotional support conversations, we conduct human evaluation by comparing the generated dialogues with and without persona traits, using the following indicators: Annotators evaluate the dialogues based on the following metrics: (1) Suggestion evaluates how effectively the supporter provided helpful advice. (2) Consistency assesses whether participants consistently maintained their roles and exhibit coherent behavior. (3) Comforting examines the supporter’s ability to provide emotional support to the seeker. (4) Identification determines which supporter delved deeper into the seeker’s situation and was more effective in identifying issues. (5) Overall assesses the overall performance about these two groups of dialogues. We recruited 10 native English speakers from undergraduate students in various disciplines, who had previously completed various annotation tasks and were experienced in the field. Each evaluator was compensated with a standard hourly wage and was fully informed about the tasks they needed to complete. They reviewed 50 randomly selected groups of instances from the generated dialogues with and without personas.
@@ 6.3 Case Study
A case study supports our point can be found in Figure 10, and the persona card and dialogue history can be found in Appendix F. Although we found that these dialogues provided the same level of direct emotional support, in the dialogue generated with persona, there is a tendency to use rhetorical questions to encourage the seeker to reflect and explore the content more deeply, while also offering suggestions more tactfully. In contrast, the dialogue generated without persona is more likely to provide direct affirmations or suggestions. In psychology, when a message is highly relevant to the recipient and they are already motivated to process it, rhetorical questions can enhance the persuasiveness of messages with weak arguments (Petty et al., 1981). Since emotional support is often regarded as a form of weak argument, incorporating rhetorical questions into a supporter’s response can make it more acceptable to the seeker, and facilitate deeper and more meaningful conversations. This observation highlights the role of personas in enhancing the quality of dialogue generation.
@@ 7 Conclusion
This analytical study highlights the significant potential of incorporating personas into LLM-generated emotional support dialogues to enhance their effectiveness and human-likeness. Our findings confirm that LLMs can infer stable straits from personas and maintain key persona characteristics while revealing shifts in traits such as emotionality and extraversion, which affect dialogue. The integration of personas not only improves the empathetic quality of responses but also influences the distribution of emotional support strategies, ensuring dialogues are more personalized and contextually appropriate. We encourage future researchers to build on our findings to develop more adaptive, effective, and robust ESC chatbots.
@@
Additionally, we employ an omniscient perspective in our data generation process, where both seekers and supporters can access the complete information. While this approach is common in previous studies (Zheng et al., 2023; zheng2024self), it doesn’t fully reflect the real-world conversation dynamics. Future researchers may improve realism by simulating emotional support scenarios with distinct information states for each role.
@@ Appendix C Correlations on Other Datasets
In this section, we introduce the correlations of HEXACO and CSI on the CAMS (Figure 6(a), 6(b) and 6(c)) and Dreaddit (Figure 7(a), 7(b) and 7(c)) dataset. The findings are same as in the Section 4.
@@
(a) Correlations between CSI and HEXACO of CAMS personas measured by gpt-40-mini.
@@
(b) Correlations between CSI and HEXACO from CAMS measured by Claude-3.5-haiku.
@@
(c) Correlations between CSI and HEXACO from CAMS measured by LLaMA-3.1-8B-Instruct.
@@
(a) Correlations between CSI and HEXACO of Dreaddit personas measured by gpt-40-mini.
@@
(b) Correlations between CSI and HEXACO from Dreaddit measured by Claude-3.5-haiku.
@@
(c) Correlations between CSI and HEXACO from Dreaddit measured by LLaMA-3.1-8B-Instruct.

### v2

@@ Abstract
Shenghan Wu1 Yimo Zhu1 Wynne Hsu1 Mong-Li Lee1 Yang Deng2 1National University of Singapore 2Singapore Management University shenghan@u.nus.edu
@@ 1 Introduction
In emotional support conversations (ESC), the supporter aims to help the seeker to reduce stress, overcome emotional issues, and promote mental well-being. Traditionally, ESC corpora have been developed through skilled crowdsourcing Rashkin et al. (2019); Liu et al. (2021), transcription of therapist sessions Liu et al. (2023); Shen et al. (2020), or by compiling emotional question-answer pairs from online platforms Sharma et al. (2020a); Sun et al. (2021); Garg et al. (2022); Lahnala et al. (2021). However, beyond the high costs, recent research Deng et al. (2024a) highlights several limitations in these methods, including privacy concerns, variability in data quality, and fabricated user needs created by crowdworkers. With the advent of large language models (LLMs), their powerful generalization abilities enable high-quality data annotation and generation based on specific instructions Tan et al. (2024); Ding et al. (2024). Consequently, more and more recent studies Zheng et al. (2023, 2024); Qiu and Lan (2024); Wu et al. (2024) investigate the use of LLMs to generate large-scale emotional support dialogue datasets across various scenarios via role-playing at lower costs. These efforts have significantly expanded the available ESC corpora.
@@
Despite the significant strides of LLMs in synthesizing ESC, a significant issue in generative AI annotations is the Lack of Human Intuition Deng et al. (2024a). Recent studies have demonstrated that effective emotional support requires careful consideration of individual differences, including personality traits, emotional states, and contextual factors (Ma et al., 2024; Ait Baha et al., 2023; Hernandez et al., 2023). This understanding has led to increased research attention on the role of persona in emotional support dialogues (Zhao et al., 2024; Cheng et al., 2022), particularly as AI systems become more prevalent in providing such support. The importance of incorporating psychological perspectives in developing empathetic AI assistants has been emphasized by researchers. Huang et al. (2023b) argues that psychological analysis of LLMs is crucial for creating more human-like and engaging interactions. While recent work has made progress in measuring LLMs’ personality characteristics using established psychological inventories (Frisch and Giulianelli, 2024; Safdari et al., 2023), there remains a gap in understanding how these persona-related aspects influence the generation of emotional support dialogues.
@@ 2.2 Emotional Support Dialogues
Early efforts on emotional support dialogues focused on collecting emotional question-answer data from online platforms (Medeiros and Bosse, 2018; Sharma et al., 2020b; Turcan and McKeown, 2019; Garg et al., 2022). These datasets laid the groundwork for understanding user emotions, but were limited to single-turn interactions. Empathetic Dialogue dataset (Rashkin et al., 2019) addressed this by introducing multi-turn dialogues, crowd-sourced to simulate diverse empathetic interactions. ESConv (Liu et al., 2021) further advanced the field by introducing emotional support strategies collected from psychological theories, enabling chatbots to use these strategies for more empathetic and contextually appropriate responses. Subsequent studies proposed using graph networks to capture global emotions causes and user intentions (Peng et al., 2022; Deng et al., 2023b), combining multiple emotional support strategies to enhance empathy (Tu et al., 2022), developing proactive dialogue systems to lead the conversation towards positive emotions Deng et al. (2023a, 2024b, 2025), and implementing emotional support strategies and scenarios using LLMs to create the ExTES dataset (Zheng et al., 2024).
@@ 2.3 Persona-Driven Emotional Support
Recent advances have integrated personas into emotional support dialogues to enhance personalization and diversity. The ESC dataset (Zhang et al., 2024) introduced personas into the dialogue generation process. Zhao et al. (2024) proposed a framework to extract personas from existing datasets for evaluation. Additionally, personas have been incorporated into chatbots to generate personalized responses (Tu et al., 2023; Ait Baha et al., 2023; Ma et al., 2024). These developments inspire our analysis of the relationship between personas, emotional support strategies, and dialogues, focusing on the extraction and use of personas in ESC.
@@ 3 Dataset Collection
In order to study the relationship between LLM-generated emotional support conversations (ESC) and persona traits, we first need to collect a set of non-synthetic ESC data with user personas as reference. Specifically, we select three existing emotional support datasets, including ESConv (Liu et al., 2021), CAMS (Garg et al., 2022), and Dreaddit (Turcan and McKeown, 2019). ESConv is a multi-turn emotional support dialogue dataset, while CAMS and Dreaddit are derived from Reddit posts discussing mental health issues.
@@ 4 Measurement of Persona Traits (RQ1)
Table 2: Correlations between CSI and HEXACO from ESConv measured by gpt-4o-mini. P_value of all metrics are less than 0.01.
@@
Table 3: Correlations between CSI and HEXACO from ESConv measured by Claude-3.5-haiku.
@@
Table 4: Correlations between CSI and HEXACO from ESConv measured by LLaMA-3.1-8B-Instruct.
@@ 4.2 Results and Discussions
Based on the responses from these inventories, we calculate the HEXACO and CSI dimension scores for each dataset. To evaluate whether LLMs can infer stable traits from persona cards in an emotional support context, we then compute the correlations between the HEXACO and CSI dimensions within each dataset. Each scale provides a score corresponding to each response, along with the dimension to which each question belongs. Based on the LLM’s responses, we calculate the scores for each dimension. Finally, we use Pearson correlation to analyze the relationships between each dimension of HEXACO and CSI. Tables 4, 4, and 4 present the correlations between HEXACO and CSI dimensions in the ESConv dataset (Liu et al., 2021) as measured by three different LLMs. Similarly, Tables 12, 12, and 12 show the correlations in the CAMS dataset (Garg et al., 2022), while Tables 15, 15, and 15 provide the correlations observed in the Dreaddit dataset (Turcan and McKeown, 2019).
@@ 5.2 Results and Discussions
The comparisons of scores of persona traits are shown in the Figure 7, 7, and 7. The results show that the original personas and the personas extracted from the generated dialogues share similar distributions across four personality traits: Honesty-Humility, Agreeableness, Conscientiousness, and Openness to Experience. This indicates that the personas maintain consistent traits in the synthetic emotional support dialogues. However, we notice that the extracted personas tend to have higher Emotionality and lower Extraversion compared to the original personas. We believe this may be because the seeker in the emotional support dialogues is dealing with emotional issues, making them more emotional and less outgoing. A similar pattern is observed in the comparison of CSI scores (see Table 21, 21, and 21 in Appendix E). This indicates that the persona traits generally maintain consistent in the synthetic emotional support dialogues with only slight shifts.
@@ 6 Impact of Persona on LLM-simulated Emotional Support Dialogues (RQ3)
Many efforts have been made on exploring how to incorporate emotional support strategies into emotional support dialogues. Zheng et al. (2024) use LLMs to introduce emotional support strategies in synthetic dialogues. Zhang et al. (2024) further involve the concept of persona into these dialogues and analyze the usage of strategies. This raises an important question: does adding a persona have an impact on the way emotional support strategies (definitions of each strategy are shown in Appendix G) are distributed in dialogues? To investigate it, we use ESConv dialogues as history and instruct LLMs to predict how a new conversation might unfold after some time. Each future dialogue is generated in two versions: with and without persona traits (personality and communication style scores). Since ESConv provides real-world context and continuations generated by LLMs can preserve the semantic distribution of dialogues (Fan et al., 2025), our approach ensures diversity and realism in emotional support scenarios.
@@
Table 5: Strategy distribution on two different groups of synthesized dialogues by gpt-4o-mini, one generated with persona traits (PT), the other without.
@@
Table 6: Strategy distribution on two different groups of synthesized dialogues by claude-haiku-3.5.
@@
Table 7: Strategy distribution on two different groups of synthesized dialogues by LLaMA-3.1-8B-Instruct.
@@ 6.1 Analysis of Emotional Support Strategies
As demonstrated in Tables 7 and 7, the distribution of emotional support strategies differs significantly between dialogues generated with and without persona traits for both gpt-4o-mini and claude-3.5-haiku. While Table 7 shows that llama-3.1-8B-instruct exhibits less pronounced differences in strategy distributions between persona and non-persona conditions, the trends remain consistent with those observed in other models. The analysis revealed that supporters in persona-enhanced dialogues demonstrated a greater tendency to validate seekers’ emotions through questioning and provided more effective emotional comfort, as evidenced by significantly higher rates of questioning, affirmation, and reassurance strategies compared to non-persona dialogues. Conversely, supporters in dialogues without persona traits emphasized problem explanation and relied more heavily on self-disclosure, while prior research (Meng and Dai, 2021) indicate that self-disclosing chatbots perform poorly when emotional support is absent and the clear boundaries are not established. The distributions of emotional support strategies generated using either personality scores or communication style scores, as shown in Table 8, demonstrate remarkable similarity. This alignment can be attributed to the strong correlation between these trait measures, as discussed in Section 4. These results provide compelling evidence that persona traits significantly influence the distribution and application of emotional support strategies in dialogues.
@@ 6.2 Human Evaluation
To intuitively reveal the impact of persona on LLM-simulated emotional support conversations, we conduct human evaluation by comparing the generated dialogues with and without persona traits, using the following indicators: Annotators evaluate the dialogues based on the following metrics: (1) Suggestion evaluates how effectively the supporter provided helpful advice. (2) Consistency assesses whether participants consistently maintained their roles and exhibit coherent behavior. (3) Comforting examines the supporter’s ability to provide emotional support to the seeker. (4) Identification determines which supporter delved deeper into the seeker’s situation and was more effective in identifying issues. (5) Overall assesses the overall performance about these two groups of dialogues. We recruited 10 native English speakers from undergraduate students in various disciplines, who had previously completed various annotation tasks and were experienced in the field. Each evaluator was compensated at a rate of approximately 20 USD per hour and was fully informed about the tasks they needed to complete. They reviewed 50 randomly selected groups of instances from the generated dialogues with and without personas.
@@ 6.3 Case Study
A case study supports our point can be found in Figure 10, and the persona card and dialogue history can be found in Appendix F. Although we found that these dialogues provided the same level of direct emotional support, in the dialogue generated with persona, there is a tendency to use rhetorical questions to encourage the seeker to reflect and explore the content more deeply, while also offering suggestions more tactfully. In contrast, the dialogue generated without persona is more likely to provide direct affirmations or suggestions. In psychology, when a message is highly relevant to the recipient and they are already motivated to process it, rhetorical questions can enhance the persuasiveness of messages with weak arguments (Petty et al., 1981). Since emotional support is often regarded as a form of weak argument (Petty and Cacioppo, 2012), incorporating rhetorical questions into a supporter’s response can make it more acceptable to the seeker, and facilitate deeper and more meaningful conversations. This observation highlights the role of personas in enhancing the quality of dialogue generation.
@@ 7 Conclusion
This analytical study highlights the potential of incorporating personas into LLM-generated emotional support dialogues to enhance effectiveness and human-likeness. Our findings confirm that LLMs can infer stable traits from personas and maintain key persona characteristics while revealing shifts in emotionality and extraversion traits that affect dialogue. Personas not only improve the empathetic quality of responses but also influence the distribution of emotional support strategies, ensuring dialogues are more personalized and contextually appropriate. We encourage future researchers to build on our findings to develop more adaptive, effective, and robust ESC chatbots.
Acknowledgments
This research was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C012).
@@ Acknowledgments
Additionally, we employ an omniscient perspective in our data generation process, where both seekers and supporters can access the complete information. While this approach is common in previous studies (Zheng et al., 2023, 2024), it doesn’t fully reflect the real-world conversation dynamics. Future researchers may improve realism by simulating emotional support scenarios with distinct information states for each role.
@@ Appendix C Correlations on Other Datasets
In this section, we introduce the correlations of HEXACO and CSI on the CAMS (Figure 12, 12 and 12) and Dreaddit (Figure 15, 15 and 15) dataset. The findings are same as in the Section 4.
@@
Table 10: Correlations between CSI and HEXACO of CAMS personas measured by gpt-40-mini.
@@
Table 11: Correlations between CSI and HEXACO from CAMS measured by Claude-3.5-haiku.
@@
Table 12: Correlations between CSI and HEXACO from CAMS measured by LLaMA-3.1-8B-Instruct.
@@
Table 13: Correlations between CSI and HEXACO of Dreaddit personas measured by gpt-40-mini.
@@
Table 14: Correlations between CSI and HEXACO from Dreaddit measured by Claude-3.5-haiku.
@@
Table 15: Correlations between CSI and HEXACO from Dreaddit measured by LLaMA-3.1-8B-Instruct.